OpenStudy (anonymous):

Julie is walking her dog around her neighborhood. She walks 1.5 miles in 45 minutes. What is her average speed? How did you find out? Julie is walking her dog around her neighborhood. She walks 1.5 miles in 45 minutes. What is her average speed? How did you find out? @Mathematics

OpenStudy (anonymous):

Average speed = distance divided by time. Typically, you will have to convert time into hours or seconds. Although, minutes is not disallowed.

OpenStudy (anonymous):

So I Have To Convert 45 Mins. Into Seconds ?

OpenStudy (anonymous):

You are better off converting into hours as people measure speed of walking as miles per hour. In any case, to convert minutes into seconds, you have to multiply the minutes by 60 as each minute is 60 seconds.

OpenStudy (anonymous):

So Multiply 45 By 60

OpenStudy (anonymous):

To convert into seconds, yes. To convert into hours, you need divide the minutes by 60 because there are 60 minutes in an hour. So, it is 45/60 = 3/4 hour.

OpenStudy (anonymous):

So It's .75?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Now, you need to find average speed as: Average speed = distance divided by time.

OpenStudy (anonymous):

Well It Says That Is Incorrect. It's Not Even An Answer On Multiple Choice

OpenStudy (anonymous):

It Says Its 2.0

OpenStudy (anonymous):

0.75 is the "hours" it took to go 1.5 miles. To get the average speed, you need to now calculate 1.5/0.75 which will exactly be equal to 2.0 miles per hour.

OpenStudy (anonymous):

Okay . So , You Had To Convert The minutes Into Hours. Then , Divide It By 60, Which Equals .75 Then divide That by 1.5 = 2.0

OpenStudy (anonymous):

Yes. So, you were given time in minutes. First convert to hours. We got 45/60 = 0.75 hours (or 3/4th hour). Now, you have to calculate miles per hour. So, 1.5 miles / 0.75 hours = 1.5 / (3/4) = 1.5 * 4 / 3 = 0.5 * 4 = 2.0 miles per hour.

OpenStudy (anonymous):

Oh , Okay. Now I Understand. Thank You